### v1

Abstract
We propose a novel method for fitting machine-learning interatomic potentials with magnetic degrees of freedom, namely, magnetic Moment Tensor Potentials (mMTP). The main feature of the methodology consists in fitting mMTP to magnetic forces (negative derivatives of energies with respect to magnetic moments) derived from spin-polarized density functional theory calculations. We test our method on the bcc Fe-Al system with different composition. Specifically, we calculate formation energies, equilibrium lattice parameter, and total cell magnetization. Our findings demonstrate a precise match between values calculated with mMTP and those obtained by DFT at zero temperature. Additionally, using molecular dynamics, we estimate the finite temperature lattice parameter and capture the cell expansion as was previously revealed in experiment. We demonstrate that mMTPs fitted to magnetic forces, increase the relaxation reliability, which is the percent of successfully relaxed structures (i.e. with almost zero forces, stresses, and magnetic moments after the optimization of geometry). Eventually, we show that the proposed methodology can provide an accurate and reliable mMTP with reduced number of computationally complex spin-polarized density functional theory calculations.
Keywords: magnetic machine-learning interatomic potentials, constrained density functional theory, molecular dynamics, magnetism.
Introduction
Machine-learning interatomic potentials (MLIPs) have become a reliable and computationally efficient tool for materials modeling. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Over the past few years a remarkable progress has been achieved in the development of magnetic MLIPs [12, 13, 14, 15, 16, 17, 18, 19, 20], whose functional form explicitly includes magnetic degrees of freedom. These models can be used to predict magnetic properties of materials such as magnetization, Curie temperature, Neel temperature, magnon dispersion, phonon spectra and elastic constants in different magnetic states.
The training set for fitting magnetic MLIPs is typically obtained from spin-polarized density functional theory (DFT) calculations. However, such calculations bear even higher computational burdening than non spin-polarized DFT due to explicit evaluation of magnetic moments of atoms. Besides, accurate training of magnetic MLIPs require a significantly larger training set due to doubling of magnetic MLIPs parameters ($6N$ against $3N$ in non-magnetic MLIPs). Therefore, reduction of the training set size for an efficient fitting of accurate magnetic MLIPs is an urgent problem.
Traditionally, non-magnetic MLIPs are trained to energies, forces and, optionally, stresses, derived from DFT and reproduce these quantity. However, with spin-polarized DFT calculations, we can obtain an additional quantity, which is a negative derivative of energy with respect to magnetic moments, known as magnetic forces. Therefore, to obtain a robust magnetic MLIPs, they can be also trained to reproduce this quantity. The opportunity was first explored by Yuan et al. [20] where the authors fitted equivariant neural network force field (NNFF) to energies, forces, and magnetic forces and accurately predicted magnon dispersion of CrI3 monolayer with the obtained model. More importantly, an accurate NNFF fitting required a relatively small training set size ($<1000$ configurations). This result demonstrates that fitting magnetic MLIPs to magnetic forces can be a possibility to obtain an accurate model using a limited set of spin-polarized DFT calculations.
In this work, we propose a novel method for fitting a magnetic MLIP, specifically, magnetic Moment Tensor Potential (mMTP) [16, 18]. The main principle consists in fitting mMTPs to magnetic forces in addition to energies, forces, and stresses. The training set is obtained using the recently developed constrained DFT (cDFT) method [21] allowing to impose hard constraints on magnetic moments and calculate magnetic forces. Additionally, we employ non-constrained DFT calculations for obtaining the configurations with the equilibrium magnetic moments (i.e., zero magnetic forces). The resulting training set includes 2632 16-atom bcc Fe-Al configurations with different concentrations of Al atoms (up to 50 %) and different random distributions of the atom types within the supercell. Using the obtained training set, we construct two ensembles of mMTPs using two methods. The first method consists in training mMTPs only on energies, forces, and stresses [16, 18]. In the second method, presented in this work, we additionally fit mMTPs to magnetic forces. Further, the two methods are compared in terms of the predictive ability of the fitted mMTPs. Specifically, we predict formation energies, lattice parameters, and total magnetic moment of the unit cell for different concentrations of Al in the Fe-Al system. Additionally, we employ mMTPs trained on magnetic forces in finite temperature molecular dynamics simulation and reproduce the cell expansion trend. Finally, we assess the reliability of the obtained models by the percent of configurations with successfully equilibrated magnetic moments (i.e. almost zero magnetic forces) and configurations with fully relaxed geometry. We note that mMTPs may have up to several thousand of parameters and thus the training set of 2632 configurations is not very large for fitting mMTPs. We show that the novel method produces systematically more accurate and reliable mMTPs on the same training set in comparison with the previous method.
The paper is organized as follows. In “Methods” section we discuss the concept of magnetic MTP and the method of fitting mMTP to magnetic forces. In “Results and Discussion” we first describe training and validation procedures of the fitted ensembles of potentials. After, we perform a quantitative analysis by comparing formation energies, lattice parameters, and total magnetic moments of the unit cell calculated by two ensembles of mMTPs and DFT. In “Conclusion” section we discuss the results of this study.
1 Methods
1.1 Magnetic Moment Tensor Potential
The mMTP is a machine-learning interatomic potential based on the non-magnetic version of MTP. The original MTP was proposed by A. Shapeev [4] and further extended to multi-component non-magnetic systems [22, 23]. This potential was further modified to take into account magnetic degrees of freedom, first for the single-component magnetic Fe system [16], and then for the multi-component magnetic Fe-Al system [18].
In mMTP, total energy of configuration is a sum of energy contributions from each local atomic environment $\mathfrak{n}_{i}$ in the system:
$E=\sum\limits_{i=1}^{n}V(\mathfrak{n}_{i}),$ (1)
where $n$ is the number of atoms in the system, $\mathfrak{n}_{i}$ denotes an atomic environment (neighborhood) of the $i$-th atom, $V(\mathfrak{n}_{i})$ is the energy contribution of a neighborhood. A neighborhood is the collection of atoms around a central atom, constrained within a sphere of a cut-off radius $R_{\rm cut}$. The neighborhood is represented by the distances $|\textbf{r}_{ij}|$ between the $i$-th atom and its neighboring atoms $j$, as well as the atomic number $z$ and magnetic moment $m$ of atoms $i$ and $j$:
$\mathfrak{n}_{i}=\{(|\textbf{r}_{ij}|,z_{i},z_{j},m_{i},m_{j}),j=1,\ldots,N_{\rm nb}^{i}\},$ (2)
where $N_{\rm nb}^{i}$ is the total number of atoms in the neighborhood. It should be noted that in this mMTP form magnetic moments are scalars and, thus, only collinear magnetism can be described.
The energy $V(\mathfrak{n}_{i})$ of a local neighborhood $\mathfrak{n}_{i}$ is linearly expanded through the basis functions $B_{\alpha}$:
$V(\mathfrak{n}_{i})=\sum\limits_{\alpha}\xi_{\alpha}B_{\alpha}(\mathfrak{n}_{i}),$ (3)
where $\xi_{\alpha}$ are linear parameters determined from fitting to configurations from the training set. The basis functions $B_{\alpha}$ are defined as all possible contractions of moment tensor descriptors $M_{\mu,\nu}$:
$M_{\mu,\nu}(n_{i})=\sum\limits_{j}f_{\mu}(|r_{ij}|,z_{i},z_{j})\textbf{r}_{ij}^{\otimes\nu},$ (4)
where $\otimes$” denotes outer product of vectors. The number of contractions is limited by the potential level, which was described in the original work [4]. The radial part of descriptors $f_{\mu}$ can be written through polynomial functions:
$f_{\mu}(|r_{ij}|,z_{i},z_{j},m_{i},m_{j})=\sum\limits_{\zeta=1}^{N_{\phi}}\sum\limits_{\beta=1}^{N_{\psi}}\sum\limits_{\gamma=1}^{N_{\psi}}c_{\mu,z_{i},z_{j}}^{\zeta,\beta,\gamma}\phi_{\zeta}(|\textbf{r}_{ij}|)\psi_{\beta}(m_{i})\psi_{\gamma}(m_{j})(R_{\rm cut}-|\textbf{r}_{ij}|)^{2},$ (5)
where $c_{\mu,z_{i},z_{j}}^{\zeta,\beta,\gamma}$ are radial parameters obtained during fitting, $\phi_{\zeta}(|\textbf{r}_{ij}|),\psi_{\beta}(m_{i}),\psi_{\gamma}(m_{j})$ are Chebyshev polynomials of the order $\zeta,\beta,\gamma$, respectively, which take values from -1 to 1, and the term $(R_{\rm cut}-|\textbf{r}_{ij}|)^{2}$ provides a smooth decay when reaching the cut-off radius. The function $\phi_{\zeta}(|\textbf{r}_{ij}|)$ is defined on the interval $(R_{\rm min},R_{\rm cut})$, where $R_{\rm min}$ denotes a minimal distance between interacting atoms. The arguments of the functions $\psi_{\beta}(m_{i})$, $\psi_{\gamma}(m_{j})$ are atomic magnetic moments on the interval $(-M_{\rm max}^{z_{i}},M_{\rm max}^{z_{i}})$, where $M_{\rm max}^{z_{i}}$ is the maximal absolute value of magnetic moment for the atomic type $z_{i}$. The number of radial coefficients is equal to $N_{\mu}\cdot N_{\phi}\cdot N_{\rm types}^{2}\cdot N_{\psi}^{2}$, where $N_{\mu}$ is the number of radial functions and $N_{\rm types}$ is the number of atomic types. Thus, the number of radial parameters increases quadratically with the number $N_{\psi}$ of magnetic basis functions. In this paper we test mMTPs with different number of magnetic basis functions.
We denote all the mMTP parameters by $\bm{\theta}=\{\xi_{\alpha},c_{\mu,z_{i},z_{j}}^{\zeta,\beta,\gamma}\}$ and the total energy of configuration by $E=E(\bm{\theta},\bm{R},M)$, where $\bm{R}=({\bm{r}}_{1},\ldots,{\bm{r}}_{n})$ are atomic positions and $M=(m_{1},\ldots,m_{n})$ are magnetic moments of atoms in a configuration. The magnetic MTP is invariant with respect to inversion of magnetic moments and the total energy of configuration calculated by mMTP is:
$E^{\rm mMTP}=E^{\rm mMTP}(\bm{\theta})=\dfrac{E(\bm{\theta},\bm{R},M)+E(\bm{\theta},\bm{R},-M)}{2}.$ (6)
From Eq. (6) we conclude that the total energy of configuration calculated with mMTP depends on the magnetic moments of the atoms explicitly. Each step of any atomistic simulation with mMTP, e.g. geometry optimization or molecular dynamics simulations, is conducted in two stages. At the first stage, we equilibrate magnetic moments (i.e. minimize energy with respect to magnetic moments) for the fixed atomic positions and lattice vectors. At the second stage, we move atoms and, optionally, change lattice vectors for the fixed magnetic moments. Thus, the correct and reliable equilibration of magnetic moments is critical for mMTP. However, this requires a correct description of magnetic forces, which is discussed in the next section.
1.2 Fitting to magnetic forces
During the equilibration of magnetic moments we calculate magnetic forces as:
$\textbf{T}=-\frac{\partial E}{\partial\textbf{m}}.$ (7)
As it was mentioned above, here we consider only collinear magnetic moments $m$ and, therefore, magnetic force is a scalar denoted by $T$.
Since the mMTP energy explicitly depends on magnetic moments of atoms, such a derivative $T$ can be calculated as a function of all the parameters $\theta$ of the potential. Thus, along with DFT energies $E^{\rm DFT}$, forces $\textbf{f}^{\rm DFT}$, and stress tensor components $\sigma^{\rm DFT}$, magnetic forces $T^{\rm DFT}$ can be used to find optimal parameters of the potential by minimizing the following objective function:
$\begin{split}\sum_{k=1}^{K}\Biggl{[}w_{e}\Bigl{(}E^{\rm mMTP}_{k}(\bm{\theta})-E_{k}^{\rm DFT}\Bigr{)}^{2}+w_{f}\sum_{i=1}^{n}\Bigl{(}\textbf{f}_{i,k}^{\rm mMTP}(\bm{\theta})-\textbf{f}_{i,k}^{\rm DFT}\Bigr{)}^{2}\\ +w_{s}\sum_{a,b=1}^{3}\Bigl{(}\sigma_{ab,k}^{\rm mMTP}(\bm{\theta})-\sigma_{ab,k}^{\rm DFT}\Bigr{)}^{2}+w_{t}\sum_{i=1}^{n}\Bigl{(}T^{\rm mMTP}_{i,k}(\bm{\theta})-T^{\rm DFT}_{i,k}\Bigr{)}^{2}\Biggr{]},\end{split}$ (8)
where $w_{e},w_{f},w_{s}$, and $w_{t}$ are non-negative weights that determine the relative importance of energies, forces, stresses, and magnetic forces during mMTP fitting. Selection of the optimal weights is discussed in the next section. We also demonstrate that minimizing the objective function (8) with the optimal weight $w_{t}$ gives much more accurate magnetic forces than in the case when $w_{t}=0$ (i.e., we do not fit mMTP to magnetic forces). Finally, we show that the method including fitting to magnetic forces is useful for the case of small training set, like the one we use in this work.
2 Results and Discussion
2.1 Training set
For constructing the training set we start from 21 configurations with fully optimized geometries. Among them are 16-atom bcc Fe-Al supercells with different concentrations of Al atoms (from 0 % to 50 %) and different distributions of the atom types within the supercell. Some examples are represented in Fig. 1. The entire training set of 2632 configurations is obtained by randomly perturbing the initial 21 configurations and using 300 K molecular dynamics simulations. Thus, about 80 % of the new configurations are the ones with random perturbation of atomic positions, lattice vectors, and magnetic moments of the initial configurations as described in [18] and the remaining configurations correspond to 300 K molecular dynamics. Further, this training set is used to test our methodology.
Figure 1: The $2\times 2\times 2$ 16-atomic bcc supercell of the Fe-Al system. Some examples of configurations with different concentration of Al atoms: Fe14Al2, Fe12Al4, and Fe8Al8.
The energies, forces, stresses, and magnetic forces for given atomic positions are calculated with DFT and the constrained DFT (cDFT) method [21] implemented in ABINIT [24, 25]. We use a 6 $\times$ 6 $\times$ 6 k-point mesh with a cut-off energy of 25 Hartree and projector augmented wave (PAW) pseudopotentials, implemented within the PBE (generalized gradient approximation) DFT functional. Importantly, about 85 % of the configurations contain non-equilibrium magnetic moments and therefore non-zero magnetic forces.
2.2 Optimal weight in the case of fitting to magnetic forces
Once we included a magnetic force optimization in the objective function (Eq. (8)), we are required to select an optimal weight for fitting to magnetic forces. To determine the optimal $w_{t}$ value we fix the rest of the weights $w_{e}=1$, $w_{f}=0.01\leavevmode\nobreak\ \mbox{\AA}^{2}$, and $w_{s}=0.001$. Such choice for energy, force, and stress weights is dictated by the previous works where these values were used to accurately fit both non-magnetic and magnetic MTP [26, 16, 18]. We further train mMTPs with different magnetic force weights in the range from $10^{-4}\leavevmode\nobreak\ \mu_{B}^{2}$ to $10^{1}\leavevmode\nobreak\ \mu_{B}^{2}$ and with $N_{\psi}=2$ (415 parameters). We use 5-fold cross-validation to verify the quality of the fitted potentials. The results are represented in Fig.  2.
Figure 2: Validation root mean square errors (RMSEs) for energies, forces, stresses, and magnetic forces with different magnetic force weights in the objective function. Potentials were trained using a 5-fold cross-validation algorithm. We provide the results with 68% confidence interval (i.e. 1-$\sigma$ interval).
From Fig. 2 we can observe a decrease of root mean square error (RMSE) for energy with the increase of magnetic force weight up to $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$. The errors for forces and stress components rise monotonously with the increase of $w_{t}$. At the same time, magnetic force errors expectedly decrease when $w_{t}$ is higher. Analyzing the obtained results we suggest 0.1 $\mu_{B}^{2}$ to be the optimal weight for magnetic forces, as it provides minimal energy errors while force and stress errors are still acceptable.
2.3 Fitting errors for different magnetic Moment Tensor Potentials
After obtaining the optimal weight for magnetic forces, we compare the accuracy of the potentials fitted on the Fe-Al training set including 2632 configurations. We take different numbers of magnetic basis functions $N_{\psi}$, namely, 2, 3, 4, and 5. We also compare two fitting strategies: without fitting to magnetic forces, i.e. $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$, and with the optimal magnetic force weight $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$. For all the potentials we choose $N_{\phi}=8$, $R_{\rm cut}=4.5\leavevmode\nobreak\ \mbox{\AA}$ , $R_{\rm min}=2.1\leavevmode\nobreak\ \mbox{\AA}$, $M_{\rm max}^{\rm Fe}=3.7\leavevmode\nobreak\ \mu_{B}$, and $M_{\rm max}^{\rm Al}=0.2\leavevmode\nobreak\ \mu_{B}$. The number of parameters for each potential can be found in Tab. 1.
Table 1: Number of parameters for the potentials with different sizes of the magnetic basis.
magnetic basis size number of parameters
2 415
3 895
4 1567
5 2431
We train an ensemble of five potentials for each number of parameters using the aforementioned training set and then average the fitting errors. After that, we calculate RMSEs between the mMTP predictions and the DFT data. The RMSEs for energy, force, stress, and magnetic force RMSEs are shown in Tab. 2 and visualized in Fig. 3.
Table 2: Root mean square errors for energies, forces, stresses, and magnetic force predicted by mMTPs on the whole training set. $N_{\psi}$ denotes the magnetic basis size and $w_{t}$ is the magnetic force weight. The results are given with 95% confidence interval (i.e. 2-$\sigma$ interval).
$N_{\psi}$ $w_{t}$ energy error force error stress error magnetic force error
($\mu_{B}^{2}$) (meV/atom) (meV/Å) (GPa) (meV/$\mu_{B}$)
2 0.0 4.70 ± 0.01 70.48 ± 0.05 0.484 ± 0.001 239.4 ± 3.5
2 0.1 3.85 ± 0.19 83.07 ± 0.91 0.617 ± 0.006 30.2 ± 0.7
3 0.0 1.98 ± 0.14 59.67 ± 0.86 0.407 ± 0.012 173.0 ± 25.8
3 0.1 2.25 ± 0.07 66.89 ± 0.36 0.479 ± 0.004 17.3 ± 0.3
4 0.0 1.51 ± 0.06 54.74 ± 1.12 0.363 ± 0.004 210.1 ± 44.8
4 0.1 1.61 ± 0.16 63.01 ± 2.61 0.450 ± 0.035 15.4 ± 0.5
5 0.0 1.51 ± 0.09 54.58 ± 1.13 0.374 ± 0.013 289.9 ± 119.4
5 0.1 1.73 ± 0.06 64.40 ± 3.95 0.469 ± 0.029 15.2 ± 0.4
Figure 3: Root mean square errors for energies, forces, stresses, and magnetic forces predicted with magnetic Moment Tensor Potentials on the whole training set. $N_{\psi}$ denotes the magnetic basis size and $w_{t}$ is the magnetic force weight.
From Tab. 2 we see that with an increase in the size of the magnetic basis up to 4, the RMSEs for energy, force, and stress decrease both for zero and non-zero magnetic force weight. This trend can be explained by the higher number of mMTP parameters when using a larger magnetic basis (see Tab. 1). Potentials with $N_{\psi}=5$ have comparable accuracy and sometimes are even less accurate, than those with $N_{\psi}=4$. Thus, in this work we focus on the magnetic basis sizes of 2, 3 and 4 in all following tests. We note that the RMSE for magnetic force decreases significantly when increasing magnetic basis size only if the weight on magnetic forces is non-zero. Hence, the quality of magnetic forces prediction does not affected by magnetic basis size if mMTPs are not fitted to magnetic forces. As it is expected, with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ the RMSE for magnetic force is significantly decreased (by a factor of ten) for each number of magnetic basis functions in comparison with the case of $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$. Moreover, fitting to magnetic forces does not significantly increase energy, force, and stress RMSEs. Thus, the mMTP fitting method including training on magnetic forces is promising in terms of accuracy and the predictive ability of the fitted mMTPs.
2.4 Correlation between DFT and magnetic Moment Tensor Potential equilibrium magnetic moments
For the additional validation of the fitted ensembles of potentials we estimate the quality of atomic magnetic moment equilibration. For this purpose we choose the configurations with the equilibrium magnetic moments from the training set, equilibrate these magnetic moments with the ensembles of mMTPs, and compare the result with DFT. The correlation between the DFT and mMTPs equilibrium Fe magnetic moments for 420 configurations are shown in Fig. 4. In the figure we observe better correlation between the DFT equilibrium magnetic moments and the ones obtained with the ensembles of mMTPs fitted to magnetic forces than with the ones without such fitting. In Tab. 3 we provide RMSEs between DFT and mMTPs equilibrium magnetic moments. As it is seen in the table, the RMSEs decrease with the increase of the number of $N_{\psi}$ for the mMTPs fitted with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ opposed to the mMTPs fitted with $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$. We note that the results are presented only for the configurations in which the magnetic moments were successfully equilibrated, i.e. when the maximal absolute magnetic force smaller than $5\cdot 10^{-3}$ meV/$\mu_{B}$ after equilibration. This quantity is referred to as equilibration reliability and is represented in Tab. 3. Indeed, all the mMTPs fitted to magnetic forces result in 100 % equilibration reliability whereas in the case when $w_{t}=0.0\leavevmode\nobreak\ \mu_{B}^{2}$ the equilibration reliability is below 100 %. A possible reason of such a result is insufficient training set size for the potentials with $N_{\psi}=3$ and $N_{\psi}=4$ that were not fitted to magnetic forces. Due to the same reason RMSE between DFT and mMTPs equilibrium magnetic moments increase with the increase of the magnetic basis.
Figure 4: Correlation between equilibrium magnetic moments of Fe calculated with mMTPs and DFT. The values of equilibrium magnetic moments obtained with the ensembles of mMTPs with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ show a better correlation with DFT than when $w_{t}=0.0\leavevmode\nobreak\ \mu_{B}^{2}$.
Table 3: RMSEs between DFT and magnetic mMTPs, equilibrium magnetic moments, and equilibration reliability obtained with different ensembles of mMTPs. We provide the values with 95% confidence interval (i.e. 2-$\sigma$ interval).
$N_{\psi}$ $w_{t}$ magnetic moment error, $\mu_{B}$ equilibration reliability, %
2 0.0 0.158 ± 0.001 96.7 ± 0.2
2 0.1 0.157 ± 0.003 100.0 ± 0.0
3 0.0 0.175 ± 0.148 98.6 ± 2.3
3 0.1 0.111 ± 0.049 100.0 ± 0.0
4 0.0 0.209 ± 0.156 98.4 ± 2.3
4 0.1 0.059 ± 0.077 100.0 ± 0.0
2.5 Formation energy, lattice parameter, and total magnetic moment obtained with different magnetic Moment Tensor Potentials
To test the predictive ability of the ensembles of mMTPs depending on the number magnetic basis functions, we calculate the formation energy, lattice parameter, and total magnetic moment of the unit cell. We start from relaxation of 21 initial configurations using the mMTPs ensembles with different number of $N_{\psi}$ fitted with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ and $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$. The errors obtained during this procedure are demonstrated in Tab. 4. We also provide the relaxation reliability, which is the percent of successfully relaxed configurations (i.e. with almost zero forces, stresses and magnetic forces). As it is seen in the table, all the 21 configurations were successfully relaxed in case when mMTPs were trained with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$. On the contrary, the relaxation reliability falls below 100 % when mMTPs are not fitted to magnetic forces. Thus, the training set including 2632 is sufficient when training to additional data such as magnetic forces. We also observe that the increase of the magnetic basis size significantly improves the predictive ability of the mMTPs. Finally, our results in Tab. 4 demonstrate that the mMTPs fitted with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ yield systematically better accuracy than the mMTPs fitted with $w_{t}=0.0\leavevmode\nobreak\ \mu_{B}^{2}$ for $N_{\psi}=3$ and $N_{\psi}=4$. Due to the above reasons we provide further results only for the ensembles of mMTPs fitted to magnetic forces.
Table 4: RMSEs between DFT and mMTPs for formation energy, equilibrium lattice parameter, total cell magnetic moment per Fe atom, and the percent of successfully relaxed configurations (relaxation reliability) of 21 initial configurations. Magnetic MTPs were fitted with different magnetic basis sizes and the magnetic force weights of 0.0 $\mu_{B}^{2}$ and 0.1 $\mu_{B}^{2}$. All the RMSEs decrease with the increase of magnetic basis size. Potentials with non-zero magnetic force weight show better accuracy compared to $w_{t}=0$ for $N_{\psi}=3$ and $N_{\psi}=4$. The results for mMTPs fitted with 0.0 $\mu_{B}^{2}$ are presented only for the successfully relaxed configurations. The values are provided with 95% confidence interval (i.e. 2-$\sigma$).
$N_{\psi}$ $w_{t}$ formation energy error lattice parameter error total magnetic moment error relaxation reliability
($\mu_{B}^{2}$) (meV/atom) (Å) ($\mu_{B}$ / $\rm N_{Fe}$) (%)
2 0.0 3.20 ± 0.03 0.0089 ± 0.0001 0.176 ± 0.001 95.2 ± 0.0
2 0.1 2.9 ± 0.4 0.0093 ± 0.0003 0.177 ± 0.001 100.0 ± 0.0
3 0.0 1.8 ± 0.8 0.0081 ± 0.0019 0.133 ± 0.071 97.1 ± 4.7
3 0.1 1.5 ± 0.1 0.0061 ± 0.0002 0.079 ± 0.027 100.0 ± 0.0
4 0.0 2.0 ± 0.5 0.0063 ± 0.0011 0.110 ± 0.051 90.5 ± 8.5
4 0.1 1.1 ± 0.3 0.0049 ± 0.0006 0.032 ± 0.003 100.0 ± 0.0
Figure 5: Formation energy of configurations with different Al concentration calculated using the ensembles of mMTP with different magnetic basis size fitted with $w_{t}=0.1\mu_{B}^{2}$ and DFT. The results are provided with a 95% confidence interval, i.e. 2-$\sigma$. All mMTPs correctly reproduce the trend of DFT formation energies.
The formation energies of 21 initial configurations calculated with different ensembles of mMTPs and DFT are shown in Fig. 5. All the ensembles of mMTPs accurately reproduce the formation energies for configurations with different concentration of Al and different distribution of atom types within the supercell. We provide further results for the configurations with the lowest formation energy for each concentration of Al in the Fe-Al compound. In Fig. 6(a) and Fig. 6(b), we present the dependence of total magnetic moment of the Fe-Al unit cell and lattice parameter on the concentration of Al. The figures reveal that values predicted with mMTPs approach the DFT-calculated values with the increase of $N_{\psi}$. Interestingly, the ensemble of mMTPs with $N_{\psi}=4$ is able to accurately reproduce the total magnetic moment of the unit cell for the Al concentration of 50 % as opposed to the ensembles with $N_{\psi}=2$ and $N_{\psi}=3$ (see Fig. 6(a)).
Figure 6: (a) Total magnetic moment of the unit cell divided by the number of Fe atoms and (b) equilibrium lattice parameter at 0 K for the configurations with minimal formation energy calculated using the ensembles of mMTPs with different magnetic basis sizes and DFT. All the mMTPs are trained with a magnetic force weight of $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$. We provide the results with the 95% confidence interval, i.e. the 2-$\sigma$ interval. The mMTP ensembles reproduce the main trend of DFT calculations for both total magnetic moments and lattice parameters, and the accuracy of mMTPs prediction improves with the increase of the number of magnetic basis functions. This improvement is clearly seen from the total magnetic moment for an Al concentration of 50 %: the potentials with $N_{\psi}=2$ give zero magnetic moment, whereas the mMTPs with $N_{\psi}=4$ correctly reproduce the DFT magnetic moment.
2.6 Molecular dynamics with magnetic Moment Tensor Potential
For testing the mMTP at finite temperature, we conduct molecular dynamics simulations in the NPT-ensemble using the LAMMPS package [27, 28]. Lattice expansion modeling was performed for a bcc supercell of 128 atoms at 300 K. We used the most accurate ensemble of mMTPs with $N_{\psi}=4$, $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ trained on the entire dataset. The obtained trend between lattice parameter and Al concentration at 300 K is depicted and compared to the experimental data [29] in Fig. 7. Additionally, we provide the lattice parameter calculated with mMTPs and DFT at zero temperature. The supercell extension trend at finite temperature is clearly observed in comparison with the zero temperature calculations. Thus, the finite temperature lattice parameter approach the experimentally measured value. However, the RMSE between the mMTPs and the experimental lattice parameters at 300 K obtained through the “cast and quenched” approach is about 0.025 Å which is larger than the RMSE between mMTPs and DFT lattice parameters of 0.005 Å obtained at zero temperature. A possible reason is the intrinsic underestimation of the lattice parameter of bcc Fe with spin-polarized PBE calculations [30] that consequently affects the underestimation of bcc Fe-Al alloys [31]. Selection of more accurate parameters for DFT calculations or the use of other exchange-correlation potential can improve the predictive ability of the model. Nevertheless, the experimental trend is correctly captured by mMTP.
Figure 7: Lattice parameters calculated at $T=0$ K and $T=300$ K using the ensemble of mMTPs fitted with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ and a $N_{\psi}=4$. Experimental points at $T=300$ K are taken from [29] and DFT calculations were obtained at $T=0$ K. The results obtained with mMTPs are provided with the 95% confidence interval, i.e., the 2-$\sigma$ interval. We observe the extension of the lattice at 300 K. The mMTPs qualitatively describe the experimentally observed expansion trend of at $T=300$ K, but quantitatively underestimate the values of the lattice parameters.
Conclusion
In this paper we developed a novel method for fitting magnetic Moment Tensor Potentials (mMTPs) [16, 18]. The methodology includes fitting to magnetic forces (negative derivatives of energies with respect to magnetic moments) in addition to energies, forces, and stresses. To test the methodology we constructed a training set containing 16-atom bcc Fe-Al supercells with different concentrations of Al and Fe using density functional theory (DFT) and cDFT calculations [21] implemented in ABINIT [24]. The resulting training set included 2632 configurations. Further, we compared two mMTPs fitting methods. In the first method used in [16, 18] we fitted mMTPs only to DFT energies, forces, and stresses. The second method included fitting to magnetic forces, obtained from DFT. First, we determined an optimal weight $w_{t}$ for fitting to magnetic forces. Next, we trained the ensembles of mMTPs with different number of magnetic basis functions using the both methods. While the novel approach did not result in an observable reduction of RMSEs for energies, forces and stresses, it significantly decreased (ten times) the RMSE for magnetic forces in comparison when mMTPs is not fitted to them. Next, we observed a better agreement between DFT-calculated equilibrium magnetic moments and that obtained with the ensemble of mMTPs fitted to magnetic forces. On the contrary, mMTPs without fitting to magnetic forces showed a higher deviation from DFT-calculated values. Additionally, the mMTPs fitted to magnetic forces gave systematically better results in predicting formation energies, lattice parameters, and unit cell magnetic moments. Moreover, we verified the impact of magnetic basis size on the accuracy of prediction of bcc Fe-Al properties. Thus, we demonstrated a higher prediction accuracy when the magnetic basis size is increased. Next, we assessed the equilibration and relaxation reliability which is the percent of configurations with successfully equilibrated magnetic moments and relaxed geometry. We demonstrated that mMTPs fitted to magnetic forces always result in 100 % reliability, while mMTPs without such fitting can not successfully equilibrate and relax all the configurations from the training set. This fact indicates a lack of the data in the training set in the case when we do not fit mMTPs to magnetic forces as opposed to the novel algorithm: all of the mMTPs were able to equilibrate and relax all the configurations. Finally, we conducted molecular dynamics simulations at $T=300$ K with the ensemble of mMTPs fitted using the novel method. We correctly captured the lattice expansion trend, observed experimentally.
The obtained results reveal that the presented methodology for mMTP fitting is a solution for obtaining an accurate mMTP keeping the training set size relatively small. The method may also be very effective in combination with active learning [32, 23] as it automatically selects the representative configurations for the training set, thus minimizing the number of computationally expensive spin-polarized DFT calculations.
CRediT authorship contribution statement
Alexey S. Kotykhov: Scripting, Calculations, Validation, Data curation, Visualization, Formal analysis, Writing - Original Draft. Konstantin Gubaev: Methodology, Software. Vadim Sotskov: Software, Formal analysis, Writing - Editing. Christian Tantardini: Scripting, Formal analysis, Writing - Editing. Max Hodapp: Resources, Calculations, Writing - Editing. Alexander V. Shapeev: Methodology, Resources, Formal analysis. Ivan S. Novikov: Methodology, Software, Formal analysis, Writing - Review & Editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work was supported by Russian Science Foundation (grant number 22-73-10206, https://rscf.ru/project/22-73-10206/). Moreover, M.H. gratefully acknowledges the financial support under the scope of the COMET program within the K2 Center "Integrated Computational Material, Process and Product Engineering (IC-MPPE)" (Project No 886385). This program is supported by the Austrian Federal Ministries for Climate Action, Environment, Energy, Mobility, Innovation and Technology (BMK) and for Labour and Economy (BMAW), represented by the Austrian Research Promotion Agency (FFG), and the federal states of Styria, Upper Austria and Tyrol. Ch.T. would like to thank Prof. Stefano Baroni and Prof. Stefano de Gironcoli at SISSA Trieste for useful discussion.